# EmojiDiff: Advanced Facial Expression Control with High Identity Preservation in Portrait Generation

### v2

@@ 5 Conclusion
We introduce EmojiDiff, an innovative solution that utilizes both portrait images and RGB expression images to generate expressive portraits. To achieve this, we design a two-stage scheme including decoupling training and fine-tuning. In the training stage, we propose ID-irrelevant Data Iteration to synthesize identity-expression decoupling data pairs, enabling the model to accept the RGB expression template without interference from expression-irrelevant features. In the fine-tuning stage, we propose ID-enhanced Contrast Alignment to further harmonize identity and expression control. This approach ensures the training process through Adaptive Noise Inversion and mitigates the impact of expression modulation on identity fidelity. Experimental results demonstrate that our approach achieves accurate and nuanced facial expression control while maintaining precise identity resemblance.
@@
In this supplementary material, we provide more metrics details in Sec. A, more implementation details in Sec. B, additional experimental results in Sec. C.
A More Metrics Details
Exp. We employ MediaPipe [30] to extract 52 facial blendshapes [16] of the expression reference and the generated image, represented as $b^{ref}$, $b^{gen}$, respectively. Each blendshape $b^{gen}_{i}$ in the [0, 1] range indicates the probability of the relevant facial action occurring (e.g., eyeBlinkLeft, mouthClose). The expression difference between reference and generated image can be formulated as:
@@ A More Metrics Details
LMS To further assess differences in the expression of critical facial regions (such as eyes, pupils, and mouth), we propose evaluating the landmark movement similarity (LMS) between the reference and generated image. First, we utilize MediaPipe [30] to detect 478 facial landmarks of the expression reference and generated image, denoted as $l^{ref}$ and $l^{gen}$, respectively. Next, we select representative landmarks to calculate the movement amplitude of key facial actions, including blinking, eye movement, and mouth opening, represented as:
@@
B More Implementation Details
B.1 Detailed Implementation
We gather images of over 10,000 people displaying various expressions from video clips and in-house face databases, removing low-quality images using LIQE [52]. Then, we employ the proposed IDI to construct cross-identity, same-expression datasets. As mentioned in Sec. 3.2.2 of the main body, facial blendshapes [16] and landmark [30] differences are utilized to filter out expression-changed data during data construction. Specifically, we calculate the Exp. and LMS metrics of the original expression image and the synthesized new image, excluding the data with Exp. $\leq$0.05 and LMS $\leq$0.18. Based on synthesized data, we train the Refined E-Adapter for SD1.5 at a resolution of 512$\times$512. On SDXL, we upscale the data to a resolution of 1024$\times$1024 and train it with random scaling following IP-Adapter [49]. The learning rates of the Adam optimizer during training and fine-tuning are set to 2$\times$ $10^{-5}$ and 5$\times$$10^{-6}$, respectively, with $\beta_{1}$ = 0.5, and $\beta_{2}$ = 0.999.
In the E-Adapter, we employ CLIP [37] as the encoder for the expression branch to extract expression signals. The image-prompt and FaceID versions of IP-Adapter [49] are utilized to initialize the projection layers of the expression and identity branches, emphasizing structure and identity features, respectively. During training and fine-tuning, the parameters of the expression branch and the projection layer of the identity branch are updated, while other parameters remain fixed.
In all experiments, $\lambda_{id}$, $\lambda_{exp}$, $R_{tmax}$, $w_{id}$, and $w_{exp}$ are consistently set to 1, 1, 600, 0.08, and 10, respectively. For expression loss $\mathcal{L}_{exp}$, landmarks of the facial contour are excluded from the loss calculation.
B.2 Derivation of Adaptive Noise Inversion
@@
where $t$, $\alpha_{t}$ represent timestep and a predefined function of $t$, respectively.  When the noise perturbation is small, the noise  $\epsilon_{\theta}(z_{t},t,C)$ predicted by U-Net approximately equals to the added noise $\epsilon$ [35, 4].  During the denoising process, we can approximately reconstruct the original latent $z_{0}$ by performing single-step sampling, denoted as:
@@
 $\hat{x}_{0}&={\rm Decode}\left(z_{0}+f(t)\ast(\epsilon-\epsilon_{\theta}(z_{t},t,C))\right),\\ f(t)&=  \begin{cases}  \sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}&\text{if }t\leq R_{tmax},\\   f(R_{tmax})&\text{if }t>R_{tmax},\end{cases} \end{aligned}$ (8)
@@ C Additional Experimental Results
C.1 More Visualization Results
In Fig. 6 of the main body, we only display a few generated images due to the page limit. In this subsection, we present additional results generated by our method on SD1.5 [38] and SDXL [36] framework. As shown in Fig. 1 and 2, our method accurately transfers the reference expression to the generated image while maintaining high identity fidelity.
C.2 Data Iteration Visualization
In Fig. 5 of the main body, we show examples illustrating the effect of IDI in modifying the identities of individuals while maintaining facial expressions. As a supplement, we provide more visualization results as depicted in Fig. 3, clearly demonstrating the effectiveness of our method.
Figure 1: More qualitative results. For the given person (leftmost column), our method generates the corresponding image based on the various expression references, evaluated on SD1.5 [38] framework.
Figure 2: More qualitative result. For the given person (leftmost column), our method generates the corresponding image based on the various expression references, evaluated on SDXL [36] framework.
Figure 3: More examples of IDI. ID-irrelevant Data Iteration (IDI) is introduced to transform expression reference images into the restruct and faceswap images, thus synthesizing high-quality data pairs with differing identities and consistent expressions.

### v3

@@ 5 Conclusion
We introduce EmojiDiff, an innovative solution that utilizes both portrait images and RGB expression images to generate expressive portraits. To achieve this, we design a two-stage scheme including decoupling training and fine-tuning. In the training stage, we propose ID-irrelevant Data Iteration to synthesize identity-expression decoupling data pairs CIEP100k, enabling the model to accept the RGB expression template without interference from expression-irrelevant features. In the fine-tuning stage, we propose ID-enhanced Contrast Alignment to further harmonize identity and expression control. This approach ensures the training process through Adaptive Noise Inversion and mitigates the impact of expression modulation on identity fidelity. Experimental results demonstrate that our approach achieves accurate and nuanced facial expression control while maintaining precise identity resemblance.
[…]
@@
In this supplementary material, we provide more experimental details in Sec. A, detailed derivation of Adaptive Noise Inversion (ANI) in Sec. B, additional experimental results in Sec. C.
A More Experimental Details
A.1 Detailed Implementation
[…]
In the E-Adapter, we employ CLIP [clip] as the encoder for the expression branch to extract expression signals. The image-prompt and FaceID versions of IP-Adapter [ye2023ipa] are utilized to initialize the projection layers of the expression and identity branches, emphasizing structure and identity features, respectively. During training and fine-tuning, the parameters of the expression branch and the projection layer of the identity branch are updated, while other parameters remain fixed.
A.2 Ablation Experiments Details
In Sec. 4.3 of the main body, we experiment with various expression controller structures. As is shown in the Fig. 1, expression embeddings are injected into the diffusion model using one of the following strategies: (a) self-attention, (b) ControlNet, (c) serial cross-attention with ID embeddings, and (d) parallel cross-attention with ID embeddings.
A.3 Metric Details
Exp. We employ MediaPipe [lugaresi2019mediapipe] to extract 52 facial blendshapes [BlendshapeV2] of the expression reference and the generated image, represented as $b^{ref}$, $b^{gen}$, respectively. Each blendshape $b^{gen}_{i}$ in the [0, 1] range indicates the probability of the relevant facial action occurring (e.g., eyeBlinkLeft, mouthClose). The expression difference between reference and generated image can be formulated as:
@@ A.3 Metric Details
LMS To further assess differences in the expression of critical facial regions (such as eyes, pupils, and mouth), we propose evaluating the landmark movement similarity (LMS) between the reference and generated image. First, we utilize MediaPipe [lugaresi2019mediapipe] to detect 478 facial landmarks of the expression reference and generated image, denoted as $l^{ref}$ and $l^{gen}$, respectively. Next, we select representative landmarks to calculate the movement amplitude of key facial actions, including blinking, eye movement, and mouth opening, represented as:
@@
Figure 1: Different expression controller structure.
Figure 2: Images generated using different prompts.
Method ID $\uparrow$ Exp. $\downarrow$ LMS $\downarrow$
DaGAN [hong2022depth] 0.202 0.135 0.569
Follow-U-Emoji [ma2024followyouremoji] 0.243 0.125 0.428
AvatarArtist [liu2025avatarartist] 0.239 0.110 0.454
SkyReels-A1 [qiu2025skyreels] 0.277 0.104 0.387
Ours 0.304 0.095 0.359
Table 1: Comparisons with post-processing animation methods.
B Derivation of Adaptive Noise Inversion
@@
where $t$, $\alpha_{t}$ represent timestep and a predefined function of $t$, respectively.  When the noise perturbation is small, the noise  $\epsilon_{\theta}(z_{t},t,C)$ predicted by U-Net approximately equals to the added noise $\epsilon$ [peng2024portraitbooth, chen2023photoverse].  During the denoising process, we can approximately reconstruct the original latent $z_{0}$ by performing single-step sampling, denoted as:
@@
[…]
@@ C Additional Experimental Results
C.1 More Quantitative Comparisons
As shown in the Tab. 1, we provide comparisons with more methods (e.g., GAN-based, animation-based, and avatar-based). Benefiting from RGB-level expression input and simultaneous processing with identity information, our method significantly surpasses other methods.
C.2 Combined with Text Prompts
As depicted in Fig. 2, we show how to combine image control and text control to generate images. Actually, prompts can be customized to meet diverse requirements while expressions are stably governed by the expression template. Moreover, an appropriate prompt will further enhance the vividness of expressions (last column).
C.3 More Visualization Results
In Fig. 5 of the main body, we only display a few generated images due to the page limit. In this subsection, we present additional results generated by our method on SD1.5 [rombach2022sd] and SDXL [podell2023sdxl] framework. As shown in Fig. 3 and 4, our method accurately transfers the reference expression to the generated image while maintaining high identity fidelity.
C.4 Data Visualization
In Fig. 4 of the main body, we show examples illustrating the effect of IDI in modifying the identities of individuals while maintaining facial expressions. As a supplement, we provide more visualization results as depicted in Fig. 5, clearly demonstrating the effectiveness of our method and data quality of CIEP100k.
Figure 3: More qualitative results. For the given person (leftmost column), our method generates the corresponding image based on the various expression references, evaluated on SD1.5 [rombach2022sd] framework.
Figure 4: More qualitative result. For the given person (leftmost column), our method generates the corresponding image based on the various expression references, evaluated on SDXL [podell2023sdxl] framework.
Figure 5: More examples of IDI and CIEP100k. ID-irrelevant Data Iteration (IDI) is introduced to transform expression reference images into the restruct and faceswap images, thus synthesizing high-quality data pairs with differing identities and consistent expressions.